CA Foundation doubt
A bag contains 5 red marbles and 3green marbles. If two marbles are drawn at random without replacement, what is the probability that both are red?
Asked by Nihla23 May 2024295 views
CA Foundation doubt
Asked by Nihla23 May 2024295 views
Hanishka · 23 May 2024
5/14
Himanshi · 23 May 2024
Both red without replacement so 5c1/8c1 ×4c1/7c1 Answer is 5/8×4/7=20/56=5/14
vrushali · 23 May 2024
2/5
Free app · 5 lakh+ learners
Read full notes, ask doubts, follow toppers and track your study hours — free on Android and iPhone.